10 liters of kerosene weighs 8 kilograms. calculate the volume of kerosene


1
Expert's answer
2021-09-21T06:33:57-0400

Given:

"V=10\:\rm L=10*10^{-3}\: m^3"

"m=8\:\rm kg"


The mass density

"\rho=\frac{m}{V}=\rm\frac{8\: kg}{10*10^{-3}\: m^3}=800\: kg/m^3"